OpenStudy (anonymous):

Calculus help please :D Find all values of x (if any) where the tangent line to the graph of the given equation is horizontal: y = −7x^2 − 4x

OpenStudy (anonymous):

remember a horizontal line a slope of zero. So you have to find the derivative and and set it equal to zero. Then take them and put it in the original forumula to fine the points y'=-14x-4 -14x-4=0 -14x=4 x=-4/14 then plus that back in to the original to get the y valu to get your y

OpenStudy (anonymous):

expanding on what I wrote x=-2/7 y=-7(-2/7)^2-4(-2/7) y=4/7 thus (-2/7,and 4/7) is is where your tangent line is parallel to the x-axis

OpenStudy (anonymous):

Actually this is quite easy to find anyways since it was a parabola anyways. So if you could fine the vertex you would have found your zero slope and would have had an answer to your tangent line problem. But yes you can do it with derivatives too